OpenStudy (anonymous):

after spending $4000 for tables and $2200 for chairs, a convention center manager finds that 20% of his original budget remains. find the amount that remains. round to the nearest dollar.

OpenStudy (anonymous):

80%= $6200 ==>10%=6200/8 =$775====> 20% =2*775 =$1550

OpenStudy (anonymous):